OpenStudy (anonymous):

Which is equal to 3[f(2x)] for f(x) = 3x^2+3/2? a.15/2 b. 81/2 c. 9x^2+9/2 d.18x^2+9

OpenStudy (e.mccormick):

What part of this is giving you trouble?

OpenStudy (anonymous):

when i try to substitute the f(x) to the 3[f(2x)] i can't get the answer

OpenStudy (e.mccormick):

You substitute 2x into f(x), then multiply through the whole thing by 3.

OpenStudy (e.mccormick):

Hmmm... and yes, if that is written right, I can see why that would not get an answer.

OpenStudy (e.mccormick):

Is it supposed to be \(3[f(2)]\)?? Rather than 2x?

OpenStudy (anonymous):

the answer is 108x^2 + 9/2?

OpenStudy (anonymous):

its 2x

OpenStudy (e.mccormick):

Yes, as written, \(108x^2 + \cfrac{9}{2}\) would be correct.

OpenStudy (muhammad_nauman_umair):

if u want this answer then u should in question (3x)^2 instead of 3(x^2)

OpenStudy (anonymous):

well there was another option to solve this problem : if the correct answer is not in the given choices, write you answer

OpenStudy (e.mccormick):

Ah, yes... I see what Muhammad_Nauman_Umair is pointing out.

OpenStudy (anonymous):

thanks for your help guys @Muhammad_Nauman_Umair and @e.mccormick

OpenStudy (e.mccormick):

\((2x)^2=4x^2\) \(3\times 4x^2=12x^2\) \(3\times 12x^2=36x^2\) Which is what I got the first time... and when you said 108 I thought I had forgotten to multiply by 3.

OpenStudy (e.mccormick):

\(36x^2 + \cfrac{9}{2}\) is what it should be.

OpenStudy (muhammad_nauman_umair):

36x^2+9/2